# RBI-ThPhys-2024-15

### v2

@@ I Motivation
Resonant di-scalar searches are currently investigated by both ATLAS and CMS ATLAS:2024ish ; CMS:2024phk , and correspond to one of the prime channels for testing the scalar sector realised in nature via triple scalar couplings. Therefore, it is of high importance that the correct theoretical description of such processes is consistently applied and up-to-date tools are used in the corresponding simulations. For example, the correct interpretation of results relies on the correct understanding of the simulated process, that might include kinematic effects in differential distributions resulting from irreducible interference terms. First experimental studies showing the possible importance of interference effects in di-Higgs production were presented by the CMS Collaboration in Ref. CMS:2024phk . Similarly, as the finite width of unstable particles is a model prediction, it cannot be set to a value chosen independently of the involved masses and couplings. Note that modifications to the di-Higgs kinematic distributions can also appear in effective field theory interpretations, see e.g. Refs. Carvalho:2015ttv ; Buchalla:2018yce ; Capozi:2019xsi ; Alasfar:2023xpc .
@@ II.1 Summary of the model
where we already imposed a $\mathbb{Z}_{2}$ symmetry that is softly broken by the vacuum expectation value (VEV) of the singlet field $S$ [^1] In fact the general phenomenology and interference effects are independent of this choice and only depend on the coupling rescaling and the total width for a chosen parameter point. Applying symmetries has some impact on the available model parameter space, mainly through theoretical constraints..
@@
[…]
@@ II.2 Constraints on the model
The parameter scan used here follows the prescription in Refs. Pruna:2013bma ; Robens:2015gla ; Robens:2016xkb ; Ilnicka:2018def . We first applied all possible theoretical and experimental constraints apart from direct search limits and signal strength measurements. This is due to the fact that most of these can be implemented directly in the scan in a straightforward way. The scan then also tests agreement with direct searches and signal strength measurements using the standalone HiggsBounds and HiggsSignals version with results up to 2022. The allowed points are then checked against the most up to date constraints using the most recent version of HiggsTools 222 Exclusion criteria within HiggsTools that are used here follow the prescription in Ref. Robens:2023pax , i.e. we use the observed and not the expected sensitivity. . The $W$-boson mass was calculated according to Refs. Lopez-Val:2014jva ; Papaefstathiou:2022oyi .
@@
Concerning the cross section values used for the determination of the rates, HiggsTools applies a factorised approach, where the production cross sections are obtained by rescaling the respective values given in the Yellow Report 4 of the Higgs working group LHCHiggsCrossSectionWorkingGroup:2016ypw , rescaled by $\sin^{2}\,\alpha$ and $\cos^{2}\alpha$ for $H$ and $h$, respectively. Note that care must be taken to distinguish the validity regimes of the respective calculations: for $m_{H}\,>\,m_{\text{top}}$ the next-to-next-to-LO (NNLO) with next-to-next-to-leading logarithmic resummation value must be taken, as it contains full top quark mass dependence up to next-to-LO order (NLO), while the next-to-NNLO calculation relies on the effective theory and is therefore viable for $m_{H}\,\leq\,m_{\text{top}}$.
Figure 1: Allowed parameter space in the real singlet extension in the $\left(m_{H},\,\sin\alpha\right)$ (left) and $\left(m_{H},\,\tan\beta\right)$ (right) planes. Maximal values for the third free parameter are shown in colour coding. In addition, on the left side we also display exclusion bounds from the $W$ mass measurement (purple), as well as several resonance searches at the LHC (colour-shaded areas). See text for details.
@@ III.1 Narrow width approximation
Let us consider a process $a\,b\,\rightarrow\,S\,\rightarrow\,c\,d$, where we now assume $S$ can be produced and decay resonantly and $a,\,b,\,c,$ and $d$ are stable initial and final state particles, respectively. In the case the $s$-channel is the only contributing diagram, the matrix element can be written via the resummation of the infinite series of self-energy insertions $\Sigma(p^{2})$ between the tree-level propagators $D(p^{2})=i/(p^{2}-m_{S}^{2})$ (see Denner:2019vbn ) as
@@
$\frac{1}{|p^{2}-m^{2}+i\,m\,\Gamma|^{2}}\,\rightarrow\,\frac{\pi}{m\,\Gamma}\,\delta\left(p^{2}-m^{2}\right)$ (18)
@@ IV Simulation of di-Higgs events
To investigate interference effects between resonant and non-resonant contributions to pair production of SM-like Higgs bosons, we simulate events of gluon fusion di-Higgs production ($gg\to hh$) both at LO and NLO in perturbative quantum chromodynamics (QCD). We perform the LO simulations with the MadGraph5_amc@nlo Alwall:2014hca event generator, using an implementation of the Two Real Singlet Extension of the SM Papaefstathiou:2020lyp with the model file from gitlabandreas . While this model contains three neutral Higgs bosons, we decouple one of these in our implementation. We denote the two non-decoupled scalar fields as $h$ and $H$, where $h$ is identified as the observed Higgs boson with a mass of $m_{h}=125\,{\,\rm GeV}$, and $H$ denotes the additional heavier Higgs boson with a variable mass $m_{H}$. The model leads to contributions to the di-Higgs amplitude from both non-resonant and resonant Feynman diagrams, including interference terms. Among the non-resonant diagrams, the box diagram, $\Box$, (Figure 2, upper-left), is independent of the Higgs trilinear couplings, $\lambda_{i}$, whereas the $s$-channel diagram, $\mathrm{S_{h}}$, depends on the $\lambda_{hhh}$ coupling (Figure 2, upper-right). The resonant $s$-channel contribution, $\mathrm{S_{H}}$, depends on the $\lambda_{Hhh}$ trilinear coupling (Figure 2, bottom).
@@
Figure 2: Feynman diagrams of the dominant contributions to di-Higgs production in the SM+singlet model at $pp$ colliders. The upper-left plot displays the so-called box diagram that does not depend Higgs trilinear couplings ($\Box$). The upper-right diagram shows the $s$-channel diagram that depends on $\lambda_{hhh}$ ($\mathrm{S_{h}}$). The resonant contribution to the di-Higgs production due to an additional additional scalar $H$ ($\mathrm{S_{H}}$) is displayed in the lower diagram.
@@
    The $\phi$ and $\eta$ direction of the $b$-quarks are smeared independently by shifting the truth-level $\phi$ and $\eta$ on an case-by-case basis by an amount determined from randomly sampling a Gaussian function with a mean of 0 and and an uncertainty of $\sigma_{\phi}=\sigma_{\eta}=0.05$, which is the typical angular resolution for $b$-jets with $p_{\mathrm{T}}\approx 100\,{\,\rm GeV}$ as estimated using simulations of the CMS detector.
@@ V.1 Illustration of interference Effects
Figure 3: The upper plots show the invariant mass $m_{hh}$ and the lower plots show the transversal momenta $p_{\mathrm{T}}$ of both final state Higgs bosons for the two parameter points. The blue line corresponds to $N_{\text{tot}}$, the red line corresponds to $N_{\mathrm{S_{H}}}$, the orange line corresponds to $N_{\text{woH}}$ and the green line is the sum $N_{\mathrm{S_{H}}}+N_{\text{woH}}$ neglecting interference effects. The event numbers $N_{i}$ corresponding to process $P_{i}$ were simulated assuming full Run-2 luminosity. The bin size is $10\,{\,\rm GeV}$.
@@
Figure 4: These plots show the interference calculated from the difference between the blue line and the green line seen in Figure 3 as the green line and a statistical error calculated as the $\sqrt{N_{\text{tot}}}$ of the $P_{\text{tot}}$ process assuming full Run-2 luminosity. The upper plots show this for the invariant mass $m_{hh}$ and the lower plots show this for the transversal momentum $p_{\mathrm{T}}$ for the two parameter points.
@@ V.2 Parameter scan
Figure 5: The different bins on the x-axis show the investigated values of $m_{H}$. The points inside the bins are ordered by the corresponding $\tan\beta$ values from left to right. The y-axis shows $\sin\alpha$, where the SM case of $\sin\alpha=0$ is omitted. Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Section II. The maximal relative interference for allowed points is around $13\%$. The colours of the points indicate the size of $\left(\Delta\sigma\right)_{\text{abs}}$ and $\left(\Delta\sigma\right)_{\text{rel}}$ for the left and right plots, respectively.
We investigated the interference effects for events with invariant mass $m_{hh}$ above or below $m_{H}$. For this we used the pylhe pylhe python package to divide our generated events into those with higher and those with lower invariant mass than the respective $m_{H}$. Accordingly, we split the cross-sections into the lower invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{<}$, and for the higher invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{>}$, respectively, as well as their absolute sum
@@
and the corresponding relative interferences $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, $\left(\Delta\sigma\right)_{\text{rel}}^{>}$ and $\left(\Delta\sigma\right)_{\text{rel}}^{\sum}$, defined accordingly by normalisation to the total cross-section over the whole mass range, equivalent to Eq. (24). The results are displayed in Figure 6. As shown in Figures 3a and especially 3b, the interference seems to have a positive value for $m_{hh}<m_{H}$ and a negative value for $m_{hh}>m_{H}$ [^7] In other beyond the SM extensions, one can also find scenarios where this behaviour is reversed, see e.g. discussion in Ref. Arco:2022lai in the context of a two-Higgs-doublet model.. Therefore, although the total interference integrated over the whole mass range might be small in some scenarios, effects in differential distributions could still be large. We therefore also provide the corresponding information on the different mass regions discussed above as well as the absolute value of deviations between naive summing of the different contributions and the total process including interference. Figure 6a shows that a higher positive mixing angle enhances both $\left(\Delta\sigma\right)_{\text{abs}}^{<}$ and $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, see Figure 6b. Furthermore, Figure 6c shows that $\left(\Delta\sigma\right)_{\text{abs}}^{>}$ has the largest negative values for large mixing angles in the $m_{H}$ region around $380\,{\,\rm GeV}$, see also Figure 6d. Figure 6e and especially Figure 6f show that some points, which appeared to have negligible interference based on the total cross-sections ($\left(\Delta\sigma\right)_{\text{abs}}$ and $\left(\Delta\sigma\right)_{\text{rel}}$) displayed in Figures 5a and 5b, actually experience two interference effects of opposite sign, which partially cancel when the total cross-section is computed. In Section VII.2 we will introduce benchmark (BM) points that exhibit this feature (benchmarks BM2 and BM3).
@@
Figure 6: Interferences for invariant di-Higgs masses $m_{hh}$ above and below the heavy scalar mass $m_{H}$. Bins on the x-axis represent different $m_{H}$ values, ordered by increasing $\tan\beta$ within each bin. The y-axis shows $\sin\alpha$, excluding $\sin\alpha=0$ (SM case). Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Section II. The colours of the points indicate the size of $\left(\Delta\sigma\right)_{\text{abs}}^{<}$, $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, $\left(\Delta\sigma\right)_{\text{abs}}^{>}$, $\left(\Delta\sigma\right)_{\text{rel}}^{>}$, $\left(\Delta\sigma\right)_{\text{abs}}^{\sum}$, and $\left(\Delta\sigma\right)_{\text{rel}}^{\sum}$ for the upper-left, upper-right, middle-left, middle-right, lower-left, and lower-right, plots, respectively.
@@
Figure 7: The different bins on the x-axis show the different values of $m_{H}$ investigated and inside the bins the points are ordered by the corresponding $\tan\beta$ values from left to right. The $\sin\alpha$ value is shown on the y-axis, where $\sin\alpha=0$ is left out, as it corresponds to the SM-case. Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Section II. The colours of the points indicate the size of $\kappa_{\lambda_{hhh}}$, $\lambda_{Hhh}$, $\Gamma_{h}$, and $\Gamma_{H}$ for the upper-left, upper-right, lower-left, and lower-right plots, respectively.

### v3

@@ I Motivation
Resonant di-scalar searches are currently investigated by both ATLAS and CMS ATLAS:2024ish ; CMS:2024phk , and correspond to one of the prime channels for testing the scalar sector realised in nature via triple scalar couplings. Therefore, it is of high importance that the correct theoretical description of such processes is consistently applied and up-to-date tools are used in the corresponding simulations. For example, the correct interpretation of results relies on the correct understanding of the simulated process, that might include kinematic effects in differential distributions resulting from irreducible interference terms. Similarly, as the finite width of unstable particles is a model prediction, it cannot be set to a value chosen independently of the involved masses and couplings. Note that modifications to the di-Higgs kinematic distributions can also appear in effective field theory interpretations, see e.g. Refs. Carvalho:2015ttv ; Buchalla:2018yce ; Capozi:2019xsi ; Alasfar:2023xpc . First studies showing the possible importance of interference effects in di-Higgs production were presented by the CMS Collaboration in Ref. CMS:2024phk .
In particular, there can be non-negligible interference effects between resonant and non-resonant diagrams that affect kinematic distributions as already discussed in Refs. Dawson:2015haa ; Carena:2018vpt ; DiMicco:2019ngk . These effects have been neglected by all experimental LHC searches for resonant di-Higgs production to date. In fact, most searches disregard the non-resonant contribution to the di-Higgs spectrum entirely i.e. it is neither taken into account as part of the signal or the background. Due to the complicated nature of these analyses, it is then not always possible to reinterpret these experimental results to set limits on realistic physics models. A careful examination of the interference effects is thus essential for exploring realistic new physics models in future LHC runs. Moreover, incorporating the non-resonant process and interference effects can even enhance the sensitivity of future searches as it allows for the exploration of deviations in kinematic distributions away from the resonance peak. This provides an additional handle for differentiating between the SM and new physics scenarios.
In this work, we therefore concentrate on the most simple model, adding an additional CP even neutral scalar to the SM, which has already been vastly explored by the experimental collaborations. We show that even for such a simple scenario, a description of the signal as a resonant peak in the invariant mass spectrum does not suffice to correctly describe the behaviour of the experimental distributions in all possible scenarios and that interference effects can also affect other distributions. We see this as a first example that interference effects need to be taken into account even for simple models, and we propose a reweighting method to incorporate such effects.
@@ II.1 Summary of the model
where we already imposed a $\mathbb{Z}_{2}$ symmetry that is spontaneously broken by the vacuum expectation value (VEV) of the singlet field $S$ [^1] In fact, for models which do not introduce new coloured states or more than one additional scalar boson, the general phenomenology and interference effects are independent of this choice and only depend on the coupling rescaling and the total width for a chosen parameter point. For the real singlet extension, applying symmetries only impacts the available model parameter space, mainly through theoretical constraints..
@@
In this work, we are interested in interference effects in pair production of the discovered SM-like Higgs boson, $h_{125}$, at a mass of $m_{125}=125\,{\,\rm GeV}$. We therefore identify the lighter state $h$ as $h_{125}$. Similarly, the doublet VEV is fixed to be $v=246\,{\,\rm GeV}$ from electroweak precision measurements. Hence, in total three free parameters remain:
@@ II.2 Constraints on the model
The parameter scan used here follows the prescription in Refs. Pruna:2013bma ; Robens:2015gla ; Robens:2016xkb ; Ilnicka:2018def . We first applied all possible theoretical and experimental constraints apart from direct search limits and signal strength measurements. This is due to the fact that most of these can be implemented directly in the scan in a straightforward way. As is standard, we include the following theoretical constraints:
  • boundedness from below of the potential,
  • limits from perturbative unitarity (PU),
  • limits from vacuum stability,
  • perturbativity of the couplings.
The above bounds impose limits on the potential couplings in the gauge-eigenbasis. In the scan we apply, we require all of these to be fulfilled, i.e. any parameter point that does not pass these constraints is no longer processed further. The above bounds have been discussed in great detail in Ref. Pruna:2013bma and have been further evolved on in Ref. Robens:2015gla , which we refer the reader to for more details. In general, one can e.g. say that PU puts an upper limit on $\tan\beta$ that is inversely proportional to the heavy scalar mass for the mass range considered here, while vacuum stability and perturbativity of the couplings lead to constraints in the $\left(\sin\alpha,\tan\beta\right)$ plane.
The scan then also tests agreement with direct searches and signal strength measurements using the standalone HiggsBounds and HiggsSignals version with results up to 2022. The allowed points are then checked against the most up to date constraints using the most recent version of HiggsTools 222 Exclusion criteria within HiggsTools that are used here follow the prescription in Ref. Robens:2023pax , i.e. we use the observed and not the expected sensitivity. . The $W$-boson mass was calculated according to Refs. Lopez-Val:2014jva ; Papaefstathiou:2022oyi .
@@
Concerning the cross section values used for the determination of the rates, HiggsTools applies a factorised approach, where the production cross sections are obtained by rescaling the respective values given in the Yellow Report 4 of the Higgs working group LHCHiggsCrossSectionWorkingGroup:2016ypw , rescaled by $\sin^{2}\,\alpha$ and $\cos^{2}\alpha$ for $H$ and $h$, respectively. Note that care must be taken to distinguish the validity regimes of the respective calculations: for $m_{H}\,>\,m_{\text{top}}$ the next-to-next-to-LO (NNLO) with next-to-next-to-leading logarithmic resummation value Harlander:2009mq ; Pak:2009dg ; Harlander:2009bw ; Harlander:2009my must be taken, as it contains full top quark mass dependence up to next-to-LO order (NLO), while the next-to-NNLO calculation Anastasiou:2014lda ; Anastasiou:2014vaa ; Anastasiou:2015vya relies on the effective theory and is therefore viable for $m_{H}\,\leq\,m_{\text{top}}$.
Figure 1: Allowed parameter space in the real singlet extension in the $\left(m_{H},\,\sin\alpha\right)$ (left) and $\left(m_{H},\,\tan\beta\right)$ (right) planes. The circular dots indicate allowed regions of the parameter space. Maximal values for the third free parameter are shown in colour coding. In addition, on the left side we also display exclusion bounds from the $W$-mass measurement in solid purple. Additionally, regions where LHC resonance searches exclude parts of the parameter space are indicated with colour-shaded areas. We note that these shaded areas indicate that the searches are sensitive to certain portions of the $\tan\beta$ range, but do not imply a full exclusion of the entire $\tan\beta$ range. See text for details.
@@ III.1 Narrow width approximation
Let us consider a process $a\,b\,\rightarrow\,S\,\rightarrow\,c\,d$, where we now assume $S$ can be produced and decay resonantly and $a,\,b,\,c,$ and $d$ are stable initial and final state particles, respectively. In the case the $s$-channel is the only contributing diagram, the matrix element can be written via the resummation of the infinite series of self-energy insertions $\Sigma(p^{2})$ between the tree-level propagators $D(p^{2})=i/(p^{2}-m_{S}^{2})$ (see Ref. Denner:2019vbn ) as
@@
$\frac{1}{|p^{2}-m^{2}+i\,m\,\Gamma|^{2}}\,\rightarrow\,\frac{\pi}{m\,\Gamma}\,\delta\left(p^{2}-m^{2}\right).$ (18)
@@ IV Simulation of di-Higgs events
To investigate interference effects between resonant and non-resonant contributions to pair production of SM-like Higgs bosons, we simulate events of gluon fusion di-Higgs production ($gg\to hh$) for 13 TeV proton-proton collisions both at LO and NLO in perturbative quantum chromodynamics (QCD). We perform the LO simulations with the MadGraph5_amc@nlo Alwall:2014hca event generator, using an implementation of the Two Real Singlet Extension of the SM Papaefstathiou:2020lyp with the model file from gitlabandreas . While this model contains three neutral Higgs bosons, we decouple one of these in our implementation. We denote the two non-decoupled scalar fields as $h$ and $H$, where $h$ is identified as the observed Higgs boson with a mass of $m_{h}=125\,{\,\rm GeV}$, and $H$ denotes the additional heavier Higgs boson with a variable mass $m_{H}$. The model leads to contributions to the di-Higgs amplitude from both non-resonant and resonant Feynman diagrams, including interference terms. Among the non-resonant diagrams, the box diagram, $\Box$, (Figure 2, upper-left), is independent of the Higgs trilinear couplings, $\lambda_{i}$, whereas the $s$-channel diagram, $\mathrm{S_{h}}$, depends on the $\lambda_{hhh}$ coupling (Figure 2, upper-right). The resonant $s$-channel contribution, $\mathrm{S_{H}}$, depends on the $\lambda_{Hhh}$ trilinear coupling (Figure 2, bottom).
@@
Figure 2: Feynman diagrams of the dominant contributions to di-Higgs production in the SM+singlet model at $pp$ colliders. The upper-left plot displays the so-called box diagram that does not depend Higgs trilinear couplings ($\Box$). The upper-right diagram shows the $s$-channel diagram that depends on $\lambda_{hhh}$ ($\mathrm{S_{h}}$). The resonant contribution to the di-Higgs production due to an additional scalar $H$ ($\mathrm{S_{H}}$) is displayed in the lower diagram.
@@
    The $\phi$ and $\eta$ direction of the $b$-quarks are smeared independently by shifting the truth-level $\phi$ and $\eta$ on an case-by-case basis by an amount determined from randomly sampling a Gaussian function with a mean of 0 and an uncertainty of $\sigma_{\phi}=\sigma_{\eta}=0.05$, which is the typical angular resolution for $b$-jets with $p_{\mathrm{T}}\approx 100\,{\,\rm GeV}$ as estimated using simulations of the CMS detector.
@@ V.1 Illustration of interference Effects
Figure 3: The upper plots show the invariant mass $m_{hh}$ and the lower plots show the transverse momenta $p_{\mathrm{T}}$ of both final state Higgs bosons for the two parameter points. The blue line corresponds to $N_{\text{tot}}$, the red line corresponds to $N_{\mathrm{S_{H}}}$, the orange line corresponds to $N_{\text{woH}}$ and the green line is the sum $N_{\mathrm{S_{H}}}+N_{\text{woH}}$ neglecting interference effects. The event numbers $N_{i}$ corresponding to process $P_{i}$ were simulated assuming full Run-2 luminosity. The bin size is $10\,{\,\rm GeV}$.
@@
Figure 4: These plots show the interference calculated from the difference between the blue line and the green line seen in Figure 3 as the green line and a statistical uncertainty, for which we used the $\sqrt{N_{\text{tot}}}$ of the $P_{\text{tot}}$ process assuming full Run-2 luminosity. The upper plots show this for the invariant mass $m_{hh}$ and the lower plots show this for the transverse momentum $p_{\mathrm{T}}$ for the two parameter points.
@@ V.2 Parameter scan
Figure 5: The different bins on the x-axis show the investigated values of $m_{H}$. The points inside the bins are ordered by the corresponding $\tan\beta$ values from left to right. The y-axis shows $\sin\alpha$, where the SM case of $\sin\alpha=0$ is omitted. Squares represent allowed points and crosses represent the points that do not pass the scan constraints described in Section II. The maximal relative interference for allowed points is around $13\%$. The colours of the points indicate the size of $\left(\Delta\sigma\right)_{\text{abs}}$ and $\left(\Delta\sigma\right)_{\text{rel}}$ for the left and right plots, respectively.
We investigated the interference effects for events with invariant mass $m_{hh}$ above or below $m_{H}$. For this we used the pylhe pylhe python package to divide our generated events into those with higher and those with lower invariant mass than the respective $m_{H}$. Accordingly, we split the cross-sections into the lower invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{<}$, and for the higher invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{>}$, respectively, and define their absolute sum as
@@
and the corresponding relative interference $\left(\Delta\sigma\right)_{\text{rel}}^{<}$ is defined by normalization to the total cross-section over the whole mass range,
$\left(\Delta\sigma\right)_{\text{rel}}^{\sum}\,\equiv\,\frac{\left(\Delta\sigma\right)_{\text{abs}}^{\sum}}{\sigma_{P_{\text{tot}}}}.$ (26)
The relative interferences $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, $\left(\Delta\sigma\right)_{\text{rel}}^{>}$, are also defined equivalent to Eq. (26). The results are displayed in Figure 6. As shown in Figures 3a and especially 3b, the interference seems to have a positive value for $m_{hh}<m_{H}$ and a negative value for $m_{hh}>m_{H}$ [^7] In other beyond the SM extensions, one can also find scenarios where this behaviour is reversed, see e.g. discussion in Ref. Arco:2022lai in the context of a two-Higgs-doublet model.. Therefore, although the total interference integrated over the whole mass range might be small in some scenarios, effects in differential distributions could still be large. We therefore also provide the corresponding information on the different mass regions discussed above as well as the absolute value of deviations between naive summing of the different contributions and the total process including interference. Figure 6a shows that a higher positive mixing angle enhances both $\left(\Delta\sigma\right)_{\text{abs}}^{<}$ and $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, see Figure 6b. Furthermore, Figure 6c shows that $\left(\Delta\sigma\right)_{\text{abs}}^{>}$ has the largest negative values for large mixing angles in the $m_{H}$ region around $380\,{\,\rm GeV}$, see also Figure 6d. Figure 6e and especially Figure 6f show that some points, which appeared to have negligible interference based on the total cross-sections ($\left(\Delta\sigma\right)_{\text{abs}}$ and $\left(\Delta\sigma\right)_{\text{rel}}$) displayed in Figures 5a and 5b, actually experience two interference effects of opposite sign, which partially cancel when the total cross-section is computed. In Section VII.2 we will introduce benchmark (BM) points that exhibit this feature (benchmarks BM2 and BM3).
@@
Figure 6: Interferences for invariant di-Higgs masses $m_{hh}$ above and below the heavy scalar mass $m_{H}$. Bins on the x-axis represent different $m_{H}$ values, ordered by increasing $\tan\beta$ within each bin. The y-axis shows $\sin\alpha$, excluding $\sin\alpha=0$ (SM case). Squares represent allowed points and crosses represent the points that do not pass the scan constraints described in Section II. The colours of the points indicate the size of $\left(\Delta\sigma\right)_{\text{abs}}^{<}$, $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, $\left(\Delta\sigma\right)_{\text{abs}}^{>}$, $\left(\Delta\sigma\right)_{\text{rel}}^{>}$, $\left(\Delta\sigma\right)_{\text{abs}}^{\sum}$, and $\left(\Delta\sigma\right)_{\text{rel}}^{\sum}$ for the upper-left, upper-right, middle-left, middle-right, lower-left, and lower-right, plots, respectively.
@@
Figure 7: The different bins on the x-axis show the different values of $m_{H}$ investigated and inside the bins the points are ordered by the corresponding $\tan\beta$ values from left to right. The $\sin\alpha$ value is shown on the y-axis, where $\sin\alpha=0$ is left out, as it corresponds to the SM-case. Squares represent allowed points and crosses represent the points that do not pass the scan constraints described in Section II. The colours of the points indicate the size of $\kappa_{\lambda_{hhh}}$, $\lambda_{Hhh}$, $\Gamma_{h}$, and $\Gamma_{H}$ for the upper-left, upper-right, lower-left, and lower-right plots, respectively.